OpenStudy (anonymous):

HELP PLEASE!=)The circumference of a circle is 50.24 centimeters. Find the area of the circle. (Use 3.14 for π.)

OpenStudy (johnweldon1993):

well what is the circumference of a circle? the formula?

OpenStudy (anonymous):

circumference,C = 2 *pi*r = 50.24 pi*r=50.24/2 = 25.12 \[area = (Pi times r)^{2}\/Pi] area = 25.12*25.12/3.14= 200.96...

OpenStudy (anonymous):

circumference is 50.24 and the forula is A=pi*radius^2

OpenStudy (johnweldon1993):

Circumference = 2*pi*r to find your radius...you divide 50.24 by 2....then divide it again by 3.14 now you have the radius.. to find the AREA...use your equation A = pi * r² and then you will have your area

OpenStudy (anonymous):

50.24 divided by 2 is25.12 = 12.56... 3.14 *12.56^2= 495.. ?

OpenStudy (johnweldon1993):

where did 12.56 come from? you divided the 50.24 by 2...good....by then divide THAT number by pi..(3.14)

OpenStudy (anonymous):

8

OpenStudy (johnweldon1993):

right...that's your radius now plug THAT in for "r" in your area formula

OpenStudy (anonymous):

12.56 was when i divided 25.12 by 2. my mistake

OpenStudy (anonymous):

200.96 =)

OpenStudy (johnweldon1993):

there you go :)

OpenStudy (anonymous):

thank you for helping me through it! :D